# What-if Analysis for Business Users: Current Practices and Future Opportunities

### v3

@@ 6. Study II: Follow-Up Task-Based Study
The first interview study informed various WIA techniques business users employed and the current difficulties they faced with them. In the task-based study, we implemented four representative WIA techniques identified in the previous study in a visual analytics prototype as a probe to gather participants’ hands-on experience and feedback on the benefits and enhancement opportunities of WIA techniques for making data-driven decisions.
@@ 6.1.1. Protocol
We conducted a 60-minute remote session with the same participants for the task-based study, including: (1) a tutorial, (2) a re-introduction to the marketing mix modeling (MMM) case and task, (3) a hands-on session with the WIA techniques, (4) open feedback, and (5) a post-study questionnaire, as shown in Figure 2.
@@
Tutorial Viewing. We provided a 15-minute pre-recorded video of the visual analytics prototype to acquaint participants with four representative WIA techniques implemented therein: (1) driver importance analysis, (2) sensitivity analysis, (3) goal-seeking analysis, and (4) constrained analysis, as shown in Figure 4. The demo used another common business case, deal closing analysis, which sales managers use to increase customer acquisition rate by analyzing prospective customer’s interaction data with the product and the organization.
@@
Re-Introduction to the MMM Business Use Case and Task. Participants were refreshed with the MMM use case and dataset to set context for the hands-on session. They were briefed on the same task of maximizing TV sales, but informed to use the representative WIA techniques. This lasted 5 minutes and participants could take notes if desired.
# Task-Based Study Sub-Tasks and Open-Feedback Questionnaire Sub-Tasks 1 Using driver importance analysis technique, what are the top three and least three drivers that influence the sales KPI? 2 Using sensitivity analysis technique, which advertising channels will you invest in in order of most to least to increase the sales KPI? 3 Using the sensitivity analysis technique, what is the sales achieved if you increase the Demand driver by 6% and the Supply driver by 3%? What can you say about the demand-supply relationship? 4 Using the sensitivity analysis technique, find the percentage increase in the Unit Price driver needed to achieve the same or more sales if the Demand and Supply drivers don’t change. 5 Using the summary analysis technique, discuss the trend observed of perturbing the non-advertising channel expenses drivers between a range of -100% to 700% and step size of 100% on the sales? Do you expect this behavior? 6 Using the summary analysis technique, discuss the trend observed of perturbing the advertising channel expenses. Which advertising channels would you invest or not invest in to increase the sales of the TV? Do you expect this behavior? 7 Using the goal seeking technique, what is the maximum TV sales you can achieve with the original boundary conditions? 8 What is the sales uplift that you can achieve if you only optimize the advertising channel drivers and the Unit Price driver? 9 How much will the sales increase if you have constraints of \$2-\$115 on the Unit Price driver? 10 Can you achieve a target sales of \$980 million? If yes, what is the driver that needs the maximum uplift? Do you think this is feasible? Open-Feedback Questionnaire 1 Were these specific tasks relatable to the ones you would try to answer to achieve your goal of increasing sales or any other decisions that you take in your own work? 2 Do you understand the key driving advertising channels behind sales? 3 Do you understand the relationship between the advertising channel expenses and sales? 4 How open would you be to using such WIA techniques for answering your own business questions? 5 What would be some other advanced analytics techniques that would be helpful for you to achieve your goal? 6 Were you able to reach a decision on investments needed to achieve maximum sales? Explain. 7 How confident are you in the decision you reached? Explain your reasoning. 8 What are your concerns about your inferred decision? Explain your reasoning. 9 What were the challenges you faced to achieve the goal of increasing the sales of the TV company? 10 Do you consider the analysis process you currently used of using the WIA efficient? Explain your reasoning.
@@
Hands-On Session. Participants accessed the prototype via an AWS-hosted application link. Based on a pilot study with 3 participants (not included in the 22), we gave each participant 25 minutes and broke down the task into 10 manageable sub-tasks, as tabulated in Table 4). They were instructed to think aloud, given a document detailing each sub-task, and received assistance as needed. Participants were additionally prompted to share how each technique could be applied to their own decision-making use cases.
@@
Open Feedback. Following the hands-on session, participants provided 10 minutes of feedback on the WIA techniques’ utility, benefits, concerns, and any additional analytics they wanted. This session also gathered insights on their trust, confidence, and perceived effectiveness of the techniques. Any missed feedback was collected using the open-feedback questionnaire (Table 4).
Post-Study Questionnaire. In the final 5 minutes, participants completed a post-study questionnaire (Anonymous, 2024c) comparing their own analyses practices with those supported by WIA. Details are in the supplementary materials (Anonymous, 2024a), and all sessions were recorded for further analysis.
@@ 6.2. Representative WIA Techniques Implemented into a Visual Analytics Prototype to Use as a Probe
We implemented four representative WIA techniques identified in the previous study within a visual analytics prototype to use as a probe in this task-based study. Although we focused on a subset to avoid overwhelming users, the chosen techniques covered a well-rounded set of viable techniques to help business users make data-driven decisions.
To ensure a comprehensive experience and enable making actionable decisions, we implemented techniques from across the pool mentioned by participants in the interview study including less popular ones. However, we made it clear that these techniques were only representative and the visual analytics prototype was only a probe to capture users’ practical and critical feedback, not an exhaustive set or tool. Throughout the study, participants were also encouraged to consider other advanced analytics techniques that could further enhance their decision-making.
We demonstrate the techniques are using the MMM business cases but they are applicable to other use cases described by business users like increasing customer retention, decreasing churn rate, and optimizing inventory. For two exemplar demonstrations, please refer to the supplementary materials (Anonymous, 2024a). We describe the representative WIA techniques in an order reflecting how business users typically approach understanding driver-KPI relationships. Nonetheless, findings from our study revealed that participants often preferred to use them in any order and iteratively.
@@
Driver Importance Analysis. This technique helped understand the relative importance of drivers in predicting the sales KPI, such as the top three advertising channels to prioritize. We implemented this technique as an ordered list of drivers (Figure 2B), accompanied by a bar chart (Figure 2A) showing the computed importance values, offering a simple to read, systematic, and insightful approach compared to approaches seen in existing tools. Using this, users quickly learnt that the top three and bottom three drivers for the sales KPI were CCI, PPI, and Supply, and Newspaper Ad Expenses, GRP of Radio, and GRP of Newspaper respectively.
Sensitivity Analysis. This technique helped observe the effects of adjusting driver values, such as increasing or decreasing expenses on different channels, on the sales KPI. This was provided in the form of coordinated views, comprising a perturbations pane (Figure 4C), a bar chart (Figure 4D), and an uplift pane (Figure 4E). The average sales observed on the original data was 96 million dollars (blue bar); upon perturbing the “supply” driver by 3% and “demand” driver by 5%, the predicted sales dynamically shows an increase by 3.63% (equivalent to 3.5 million dollars) (green texts), making the overall predicted sales 100 million dollars (yellow bar).
@@
Summary analysis (Figure 4G) allowed users to view multiple sensitivity scenarios at once, that is aided them in comprehending the sales KPI trend for when individual driver values underwent perturbation across a specified range of bounds (Figure 4F). A small multiples chart was used, with each multiple displaying the trend of the sales KPI impact when each advertising channel expense was varied from -100% (lower bound) to 500% (upper bound), with a step size of 100%, revealing significant sales increases for SMS and Radio channels, a slight rise for the Internet channel, and declines for TV and Newspaper channels.
@@
Goal-Seeking Analysis. The technique helped users learn predictions of the required driver values to achieve an optimum or user-desired aggregate KPI value, like to directly identify the actions needed to maximize the sales KPI. We implement it iteractively (Figure 4H), where users chose their optimization goal (maximize, minimize, or set a target goal), observed the pre-specified constraints on all drivers to initially be optimized between 0% to 300%, and ran the analysis. On maximizing the sales KPI goal, a prediction of 549.07 million dollars sales could be achieved (470.11% uplift) along with the driver values that enabled this prediction.
@@
Constrained Analysis. This technique incorporated user-defined business constraints (e.g., boundary, equality or inequality conditions) on one or more drivers and run goal-seeking analysis to get driver values satisfying these constraints. For instance, when using the previous goal-seeking analysis technique, participants observed that in order to maximize sales, the “unit price” driver needed to increase by 297.81% or 74000 dollars, which customers would not practically pay. In response, they could restrict the “unit price” driver to a 10% to 15% change in the constraints pane (Figure 4I) which reduced the sales prediction from 549.07 million to 277.84 million dollars, but with considering “unit price” driver prediction to be 14.08% within constraints, and yet showing a significant uplift of 188.49% while predicting the “unit price” to be 14.08% (Figure 4J) within given limits.
Each technique also included accuracy ($R^{2}$ measure) or error scores (error rate) of the underlying prediction models, with the MMM use case showing a fair accuracy of 0.82 (Figure 4K).
@@ 7. Findings of the Follow-Up Task-Based Study
We present our findings by first reporting on participant’s experiences of using the representative WIA techniques for the MMM business use case. Then, we discuss the potential benefits and opportunities for enhancement they identified regarding the use of advanced analytics for making data-driven decisions.
@@ 7.1. Insights From Using Representative WIA Techniques for the MMM Business Use Case
Feedback from both the open-feedback session and post-study questionnaire showed unanimous agreement that WIA techniques implemented in the visual analytics prototype as a probe were “helpful” (P2, P10, P12, P17, P20), “interesting” (P3, P8), “powerful” (P13, P16), and quick to “set [them] in the right direction” (P7, P17) for decision-making to maximize the sales KPI. All participants comfortably completed all 10 sub-tasks within the allotted time, finding them simple and easy to use, and made similar decisions, such as increasing the unit price of TVs, boosting SMS advertising spending, and cutting newspaper ad spending. They also mentioned that using WIA techniques was “better than sitting looking at the data” (P9) that they did in their existing workflows. All but one participant were excited to apply the techniques to their own business use cases, to “play around with own data [since it would be] super helpful when setting goals for the quarter” (P3) and learn data behaviors like “how it can relate to their business” (P6), “how it would be useful for drilled-down segments [subsets of their data]” (P4). The remainder participant, an operations manager, found the techniques to be “okay and wanted [them] to be more dummy-proof” (P15) suggesting their utility, but also the need for more greater customization towards their domain.
As for the usefulness of the techniques, participants collectively ranked them all as useful (Figure 5), with sensitivity analysis rated highest (most saturated stack), followed by constrained analysis, goal-seeking analysis, and driver importance analysis (least saturated stack). The goal-seeking technique required the most guidance (6 out of 22 participants), despite aligning well with their decision-making needs, highlighting the importance for explicit guidance, demonstrations, and examples for their practical use. While participants also initially struggled with interpreting the small multiples chart expressing summary analysis, they found it valuable for reducing cognitive load in experimenting with multiple what-if scenarios. However, they stressed the need for simpler visuals when presenting such advanced analyses to executive stakeholders.
Further, participants expressed a keen interest in additional advanced analytics techniques that they also mentioned in the first interview study. They sought automated features with robust drill-down capabilities, enabling segmentation analysis to help “how different geographical locations are doing?” (P14), “realize how the data behaves for different verticals of customers?” (P18), or “how to leverage existing customers and products doing well for those not performing?” (P21) They also wanted techniques to uncover correlations between drivers and KPIs and facilitate easier scenario comparisons. Participants requested enhancements to the WIA techniques, like seamlessly integrating results from one technique into another (e.g., plugging in the predictions from goal-seeking analysis back into sensitivity analysis (P21)) for continuing the analysis, and the ability to save analysis results for future comparisons (P1, P8, P14).
@@ 7.2.1. Helps Accelerate Decision-Making
Participants appreciated the ability of WIA to provide a “holistic view” (P13) of driver-KPI relationships, unlike the currently used manual approaches of attaining it using spreadsheet-like tools. More importantly, participants were happy that they could quickly observe the predictions for many scenarios encompassing various combination of drivers all together without the need for external expertise from data analysts or teams. Consequently, all participants shared that WIA would speed up their decision-making and “increase [their] efficiency” (P6). Participants were also pleasantly surprised by the capabilities of the representative techniques since they may have seen “something similar but not anywhere close to these [representative techniques]… to customize them or make your own if you want anything like these [representative techniques]…” (P11). As such, participants emphasized the transformative potential of incorporating advanced analytics in streamlining and expediting their decision-making.
@@ 7.2.2. Increases Confidence in Decisions
Participants articulated that WIA helped them move away from from guesswork and trial-and-error approaches, thereby bolstering their confidence in their decision-making (Elhamdadi et al., 2022), as it narrows down the vast possibilities of hypotheses to those observed in the data. One participant P1, for instance, happily shared how WIA can “help understand where to put more effort!” Moreover, participants emphasized that leveraging advanced analytics techniques, particularly interactively as demonstrated in our study, could help them think and test out a lot more corner cases and scenarios that they might miss in traditional manual analysis. Further, participants shared how they could play with many more drivers to observe their consequences on the business KPI and be able to understand the market, product, and customer behavior more widely and better. All these reasons combined led the participants to excitedly share how they can build a strong narrative or story supporting their decision actions. A participant P10, for example, expressed how WIA will enable “everyone in the company [to] be on the same page about a decision… no one’s decision will be overwritten”. Therefore, participants showed promise in WIA enabling them to be confident in their decisions and ready when communicating with the stakeholders.

### v4

@@ 6. Study II: Follow-Up Task-Based Study
The first interview study informed us of the workflows business professionals use to make data-driven decisions, the various WIA techniques they employ, how they currently execute them, the difficulties they face, and the techniques they desire to have. Building on these findings, in the task-based study, we wanted to learn how the participants perceived and interacted with advanced analytics techniques. To achieve this, we implemented four representative WIA techniques identified in the previous study within an interactive visual analytics prototype. This served as a probe to gather participants’ hands-on experience and feedback on using WIA techniques for making data-driven decisions. This approach aimed to address the challenges and limitations of participants’ existing workflows and tools by providing a more intuitive and streamlined experience.
@@ 6.1.1. Protocol
We conducted a 60-minute remote session with the same participants for the task-based study with the following protocol and as shown in Figure 1.
@@
Tutorial Viewing. We provided a 15-minute pre-recorded video of the visual analytics prototype to acquaint participants with the implemented WIA techniques. The demo used another common business case, deal closing analysis, which sales managers use to increase customer acquisition rate by analyzing prospective customer’s interaction data with the product and the organization.
@@
Re-Introduction to the MMM Business Use Case and Task. Participants were reintroduced to the MMM use case and dataset to establish the context for the hands-on session. They were then briefed on the same task of maximizing TV sales, but this time with instructions to apply representative WIA techniques using the prototype. This lasted 5 minutes and participants could take notes if desired.
@@
Hands-On Session. Participants accessed the prototype via an AWS-hosted application link. Based on a pilot study with 3 participants (not included in the 22), we gave each participant 25 minutes and broke down the task into 10 manageable sub-tasks, as tabulated in Table 3). They were instructed to think aloud, given a document detailing each sub-task, and received assistance as needed. Participants were additionally prompted to share how each technique could be applied to their own decision-making use cases. The coder took notes on interesting participant interactions (e.g., switching between different functionalities), clarify questions asked, or feedback on using the functionalities (e.g., goal-seeking analysis needed multiple iterations to understand).
@@
Open Feedback. Following the hands-on session, participants provided 10 minutes of feedback on the WIA techniques’ utility, benefits, concerns, and any additional analytics they wanted. This session also gathered insights on their trust, confidence, and perceived effectiveness of the techniques. The coder also took notes of the feedback to ensure that any missed feedback from the open-feedback questionnaire (Table 3) was collected.
Post-Study Questionnaire. In the final 5 minutes, participants completed a post-study questionnaire (Anonymous, 2024c). This questionnaire included Likert scale-based questions to capture participants’ thoughts on the usefulness and usability of the functionalities, as well as open questions to gather quantitative ranking of functionalities based on usefulness, concerns with functionalities, and any additional functionalities they would like. Details are in the supplementary materials (Anonymous, 2024a), and all sessions were recorded for further analysis.
@@ 6.2. Representative WIA Techniques Implemented into a Visual Analytics Prototype to Use as a Probe
We chose four WIA techniques from the seven techniques identified in the previous study and implemented them within an interactive visual analytics prototype to use as a probe in this task-based study. We did not implement all seven techniques to avoid overwhelming the participants. The four techniques were driver importance analysis, sensitivity analysis, goal-seeking analysis, and constraint analysis. They were considered representative because they either directly encompass other WIA techniques or serve as important building blocks for more complex WIA techniques. For example, segmentation analysis, which involves filtering and drilling down into specific data groups, can be viewed as a preliminary step to applying techniques like driver importance or sensitivity analysis on data subsets. Similarly, correlation analysis, which focuses on understanding relationships between drivers or between drivers and KPI, is inherently part of driver importance analysis, as it seeks to identify and quantify the impact of drivers on KPI and between each other. Outlier analysis can also be integrated within these techniques by first identifying anomalies and then applying sensitivity or constraint analysis to understand and manage their impact.
Participants were informed that the visual analytics prototype was only a probe to capture professionals’ practical and critical feedback, not an exhaustive set or tool. Throughout the study, participants were also encouraged to consider other advanced analytics techniques that could further enhance their decision-making. We also informed the participants that the techniques were applicable to other use cases beyond the MMM scenario, like increasing customer retention, decreasing churn rate, and optimizing inventory. We described the representative WIA techniques in an order reflecting how business professionals typically approach understanding driver-KPI relationships. Findings from our study revealed that participants often preferred to use them in any order and iteratively.
Driver Importance Analysis. This technique helped understand the relative importance of drivers in predicting the sales KPI, such as the top three advertising channels to prioritize. We implemented this technique as an ordered list of drivers (Figure 3B), accompanied by a bar chart (Figure 3A) showing the computed importance values, offering a simple to read, systematic, and insightful approach compared to approaches seen in existing tools. Using this, professionals quickly learned that the top and bottom three drivers influencing the sales KPI were CCI, PPI, and Supply, and Newspaper Ad Expenses, GRP of Radio, and GRP of Newspaper, respectively. Additionally, participants could interactively select a subset of drivers they were interested in to assess the importance of those specific drivers in relation to the KPI.
Sensitivity Analysis. This technique helped observe the effects of adjusting driver values, such as increasing or decreasing expenses on different channels, on the sales KPI. This was provided in the form of coordinated views, comprising a perturbations pane (Figure 3C), a bar chart (Figure 3D), and an uplift pane (Figure 3E). Participants could input a numeric or percentage change to adjust one or more drivers, which automatically updated the bar chart and uplift panel to reflect the corresponding KPI output prediction. The average sales observed on the original data was 96 million dollars (blue bar); upon perturbing the “supply” driver by 3% and “demand” driver by 5%, the predicted sales dynamically shows an increase by 3.63% (equivalent to 3.5 million dollars) (green texts), making the overall predicted sales 100 million dollars (yellow bar).
@@
[…]
[Interface of the four representative WIA techniques implemented in a visual analytics prototype.]Interface showing the four representative WIA techniques implemented in a visual analytics prototype, used as a probe to capture participants’ practical use of WIA techniques for making decisions in the MMM business use case during the task-based study: (1) driver importance analysis bar chart (A) with toggle-able drivers (B), (2) sensitivity analysis perturbation pane (C), bar chart (D), and uplift pane (E), along with its variant of summary analysis (G) with bounding pant (F), (3) goal-seeking analysis prediction (H), (4) constrained analysis options and constraints pane (I), and prediction (J), and the $R2$ accuracy score pane (K).
@@
Summary analysis (Figure 3G) allowed professionals to view multiple sensitivity scenarios at once, aiding them in comprehending the sales KPI trend for when individual driver values underwent perturbation across a specified range of bounds (Figure 3F). The perturbation range could be added as numeric values or percentage change. Small multiples chart was used, with each multiple displaying the sales KPI trend when each advertising channel expense was varied between -100% (lower bound) to 500% (upper bound), with a step size of 100%, revealing significant sales increases for SMS and Radio channels, a slight rise for the Internet channel, and declines for TV and Newspaper channels. Participants could hover over the small multiples to view a tooltip displaying the original KPI value alongside the predicted KPI value after the perturbation.
@@
Goal-Seeking Analysis. The technique helped professionals learn predictions of the required driver values to achieve an optimum or user-desired aggregate KPI value. For example, they can directly identify the actions needed to maximize the sales KPI. We implement it interactively (Figure 3H), where participants chose their optimization goal (maximize, minimize, or set a target goal), observed the pre-specified constraints on all drivers to initially be optimized between 0% to 300%, and ran the analysis. On maximizing the sales KPI goal, a prediction of 549.07 million dollars sales could be achieved (470.11% uplift) along with the driver values shown in a table having average original, predicted, and change values that enabled this prediction.
@@
Constrained Analysis. This technique incorporated user-defined business constraints (e.g., boundary, equality or inequality conditions) on one or more drivers and run goal-seeking analysis to get driver values satisfying these constraints. For instance, when using the previous goal-seeking analysis technique, participants observed that in order to maximize sales, the “unit price” driver needed to increase by 297.81% or 74000 dollars, which customers would not practically pay. In response, they could restrict the “unit price” driver to increase between only 10% to 15% by specifying it in the constraints pane (Figure 3I). Maximizing sales again reduced the sales prediction from 549.07 million to 277.84 million dollars, but ensured constraining the “unit price” driver within the specified limits to 14.08%, yet showing a significant uplift of 188.49% (Figure 3J).
The performance of the underlying prediction model was presented, either as accuracy (measured by the $R^{2}$ score) or error rate. For example, the MMM use case demonstrated a fair performance with an $R^{2}$ score of 0.82 (Figure 3K).
@@ 7. Findings of the Follow-Up Task-Based Study
We present our findings by first reporting the participants’ experiences of using the representative WIA techniques for the MMM business use case. Then, we discuss the potential benefits and opportunities for enhancement that were identified for the use of advanced analytics for making data-driven decisions.
@@ 7.1. Insights From Using Representative WIA Techniques for the MMM Business Use Case
Feedback from both the open-feedback session and post-study questionnaire showed unanimous agreement that WIA techniques implemented in the visual analytics prototype probe were “helpful” (P2, P10, P12, P17, P20), “interesting” (P3, P8), “powerful” (P13, P16), and quick to “set [business professionals] in the right direction” (P7, P17) for making decisions to maximize the sales KPI. All participants comfortably completed all 10 sub-tasks within the allotted time, finding them simple and easy to use, and made similar decisions, such as increasing the unit price of TVs, boosting SMS advertising spending, and cutting newspaper ad spending. They also mentioned that using WIA techniques was “better than sitting looking at the data” (P9) that they did in their existing workflows. All but one participant were excited to apply the techniques to their own business use cases, to “play around with own data [since it would be] super helpful when setting goals for the quarter” (P3) and learn data behaviors like “how it can relate to their business” (P6) and “how it would be useful for drilled-down segments [subsets of their data]” (P4). The remainder participant, an operations manager, found the techniques to be “okay and wanted [them] to be more dummy-proof” (P15) suggesting their utility, but with the need for greater customization towards their domain.
Participants collectively acknowledged the usefulness of all the techniques. Figure 4 illustrates the relative rankings of the techniques. Sensitivity analysis was rated the most useful, followed by constrained analysis, goal-seeking analysis, and driver importance analysis. The goal-seeking technique required the most guidance (6 out of 22 participants), despite aligning well with their decision-making needs, highlighting the importance for explicit guidance, demonstrations, and examples for their practical use. While participants also initially struggled with interpreting the small multiples chart expressing summary analysis, they found it valuable for reducing cognitive load in experimenting with multiple what-if scenarios. However, they stressed the need for simpler visuals when presenting such advanced analyses to executive stakeholders.
@@
[…]
[Ratings of the what-if analysis techniques they used in the prototype probe.]Ratings of the what-if analysis techniques from most to least useful. Participants found sensitivity analysis to be the most useful technique.
Further, participants expressed a keen interest in additional advanced analytics techniques that they also mentioned in the first interview study. They sought automated features with robust drill-down capabilities, enabling segmentation analysis to help answer “how different geographical locations are doing?” (P14), “realize how the data behaves for different verticals of customers?” (P18), or “how to leverage existing customers and products doing well for those not performing?” (P21) They also wanted techniques to uncover correlations between drivers and KPIs and facilitate easier scenario comparisons. Participants requested enhancements to currently provided WIA techniques too, like seamlessly integrating results from one technique into another (e.g., plugging in the predictions from goal-seeking analysis back into sensitivity analysis (P21)) for continuing the analysis, and the ability to save analysis results for comparisons (P1, P8, P14).
@@ 7.2.1. Helps Accelerate Decision-Making
Participants valued the ability of WIA to provide a “holistic view” (P13) of driver-KPI relationships, a stark contrast to the current manual, spreadsheet-based methods they currently rely on. More importantly, they appreciated how WIA enabled them to quickly observe predictions across multiple scenarios, incorporating various driver combinations-all without requiring external expertise from data analysts or specialized teams. Consequently, all participants agreed that WIA would accelerate their decision-making and “increase [their] efficiency” (P6). Many were also pleasantly surprised by the capabilities of the representative techniques, noting that while they had encountered “something similar, [it was] not anywhere close to these” (P11). In previous experiences, they often had to customize or build their own tools to achieve similar functionality. This underscored participants’ recognition of WIA’s transformative potential in streamlining and enhancing decision-making through advanced analytics.
@@ 7.2.2. Increases Confidence in Decisions
Participants articulated that WIA helped them move away from guesswork and trial-and-error approaches, thereby bolstering their confidence in their decision-making (Elhamdadi et al., 2022), as it narrows down the vast possibilities of hypotheses to those observed in the data. One participant P1, for instance, happily shared how WIA can “help understand where to put more effort!” Moreover, participants emphasized that leveraging advanced analytics techniques, particularly interactively as demonstrated in our study, could help them think and test out a lot more corner cases and scenarios that they might miss in traditional manual analysis. Further, participants shared how they could play with many more drivers to observe their consequences on the business KPI and be able to understand the market, product, and customer behavior more widely and better. All these reasons combined led the participants to excitedly share how they can build a strong narrative or story to support their decision actions. A participant P10, for example, expressed how WIA will enable “everyone in the company [to] be on the same page about a decision… no one’s decision will be overwritten”. Therefore, participants showed promise in WIA enabling them to be confident in their decisions when communicating with the stakeholders.